OpenStudy (anonymous):

A car starts from rest and travels for 5.0 s with a uniform deceleration of 3.0 m/s^2. -what is the final veloecity of the car(answer in m/s) -how does the car travel in this time interval?

OpenStudy (anonymous):

a=v(final)-v(initial) / t. -3m/s^2=v(final)-0/5 v(final)=-15m/s The velocity of the car decreases

OpenStudy (anonymous):

how far does it travel in this time interval?

OpenStudy (anonymous):

v=d/t v*t=d 15m/s*5s=75m